OpenStudy (anonymous):

The inflation rate of the U.S. dollar is 3.7 percent. What this means is that every year, prices increase by 3.7 percent. If a quart of oil cost $1.69 five years ago, what does it cost now?

OpenStudy (anonymous):

2.03

OpenStudy (anonymous):

multiply .037 by the original value, then add the product to the original sum. then repeat 4 more times

OpenStudy (anonymous):

hmmmm I would use the folllowing equation 1.69(1.037)^5

OpenStudy (anonymous):

169 * (1.037) to the power of 5 i guess i dont know how to add the 5 on top by the the parenthesis

OpenStudy (anonymous):

So basically we start off with 1.69 then it inflates by 3.7% So that means we have 1.69+1.69(.037)=1.69(1+.037)=1.69(1.037) And then basically we keep multiplying by 1.037 every year